OpenStudy (anonymous):

Which of the following is not a Pythagorean triple? A. 28, 45, 53 B. 16, 63, 65 C. 13, 84, 85 D. 11, 61, 62

Vocaloid (vocaloid):

a pythagorean triple will satisfy the equation a^2 + b^2 = c^2 for each answer choice, plug in the values for a, b, and c, and if the equation is satisfied, then it is a pythagorean triple

OpenStudy (anonymous):

what do you mean satisfied?

Vocaloid (vocaloid):

"satisfied" means the equation is true (both sides of the equal sign are equal)

OpenStudy (anonymous):

im confused

Vocaloid (vocaloid):

for example, we can look at answer choice A and let a = 28, b = 45, and c = 53. if a^2 + b^2 = c^2 then the equation is satisfied

OpenStudy (anonymous):

784 +2025 = 7809

Vocaloid (vocaloid):

for answer choice A, a = 28 b = 45 c = 53 a^2 + b^2 = c^2 28^2 + 45^2 = 53^2 784 + 2025 = 2809 2809 = 2809 since both sides of the equal sign are equal, the equation is satisfied, and this is a Pythagorean triple (meaning, A is not the answer) keep going until you find the answer choice that ISN'T satisfied

OpenStudy (anonymous):

so the answer would be C

Vocaloid (vocaloid):

nope, check your math again

OpenStudy (anonymous):

13, 84, 85 13 ^2 = 169 84^2 = 7056 7225

Vocaloid (vocaloid):

7059 + 169 = 7225 7225 = 7225 the equation is satisfied, so C is not our answer... please keep going with the other choices...

OpenStudy (anonymous):

D. 11, 61, 62 11 x11 = 121 61 x 61 = 3721 3844 = 63x63

Vocaloid (vocaloid):

right, D is our answer because 121 + 3721 is NOT equal to 3844

OpenStudy (anonymous):

thank you